# Upsampling DINOv2 features for unsupervised vision tasks and weakly supervised materials segmentation

### v1

@@ Acknowledgements
S8 Weakly supervised segmentation
Figure S11: Further examples of weakly supervised segmentation of materials. The first is a slice from an X-ray Computed Tomography (XCT) of glutamic acid64, displaying two crystal polymorphs: the larger, blocky $\alpha$-polymorph and the smaller, needle-like $\beta$-polymorph64. The classical features are unable to distinguish between the two phases, struggling to delineate between edges of a large $\alpha$ block and the $\beta$ needles, whilst the DINOv2 features succeed but are blurred. The combination of the two produces a good segmentation of both phases - this ability to classify using structural information (rather than focussing on pixel-value) is useful in distinguishing between polymorphs for characterization techniques that have elemental contrast (like XCT or SEM), as the polymorphs will have similar pixel-values. The second micrograph is a Scanning Electron Microscopy (SEM) image of a nickel-based superalloy59 with three phases: a large, connected matrix phase, secondary precipitates (large blobs) and tertiary precipitates (small blobs). Again, we find that the hybrid approach produces a better segmentation, especially of the third phase.
[…]
@@
Figure S13: Example segmentations from applying classifiers trained on various feature-sets to unlabelled examples in the T-cell dataset. We note the upsampled ViT features produce segmentations that much more closely align with the ground truth, able to ignore other cluttered cells, handle the presence of multiple nuclei and varying exposure. These predictions have had a CRF applied.
Figure S14: The same predictions as in Figure S13 without a CRF applied.
S9 Transformations
[…]
@@
Table S2: Performance on the cell TEM dataset for ViT features (stride 4, no CRF) without and with shift transforms applied - the performance gain is modest for DINOv2 and more marked for DINO.
@@
Table S3: Performance on DUTs foreground object segmentation as a function of transforms applied

### v2

@@ Acknowledgements
S6 Weakly supervised segmentation
S6.1 T-cell case study
We show the labels we used to train the classifiers for Section 4.4 in Figure S7. We chose cells that covered the range of variation in the dataset (size, lack of nucleus, exposure, presence of multiple cells). We found it useful to label across the space, and to focus on boundaries between the classes.
Figures S8 and Figures S9 show predictions of the classifier trained with the labels in Figure S7 on unlabelled cells in the dataset. We see greatly improved performance when using the upsampled ViT features relative to the classical features, regardless of whether CRF post-processing was used (Figure S8) or not (Figure S9).
We include FeatUp and bilinear upsampling. Bilinear upsampling performs well, especially relative to its cost. However, it struggles to capture the fine features on the edges of cells (i.e, which are smaller than the patch size of 14px), which is expected. These cells are large, so capturing fine detail is not necessary to achieve a high mIoU (in Table 4) However, in other materials, small features are highly important, like inclusions in metals or cracks in batteries, and the coarseness of the bilinear upsampling may then become a larger problem - we see this in Figure S10.
@@
Figure S8: Example segmentations from applying classifiers trained on various feature-sets to unlabelled examples in the T-cell dataset. We note the upsampled ViT features produce segmentations that much more closely align with the ground truth, able to ignore other cluttered cells, handle the presence of multiple nuclei and varying exposure. These predictions have had a CRF applied. We have abbreviated ‘DINO’ to ‘D’ i.e, ‘D-S-8’ is ‘DINO-S-8’.
Figure S9: The same predictions as in Figure S8 without a CRF applied. Again we have abbreviated ‘DINO’ to ‘D’ i.e, ‘D-S-8’ is ‘DINO-S-8’.
S6.2 Upsampler ablation
Figure S10 shows weakly supervised segmentation of the micrographs in Figure 6 using different upsampling techniques. Bilinear upsampling suffers from clear artefacting, FeatUp and our approach are better, but still display some blurring. This motivates the hybrid approach, (Figure 6), which combines the classical and upsampled ViT features to achieve a good balance of semantic and colour information.
Figure S10: Weakly supervised micrograph segmentation using different upsampled DINOv2 featuresets: bilinear upsampling, FeatUp, and our approach (stride 4, shifts 1,2 and flip transforms). We see that bilinear upsampling causes artefacts, and FeatUp, whilst better, displays some blurring (i.e, between particles).
Figure S11: Weakly supervised micrograph segmentation using different deep feature upsamplers in the ‘hybrid’ scheme (i.e, FeatUp deep features + classical or DINOv2-S-14 stride 4, shift 1,2 + classical). We see that our approach improves segmentation of fine features, like some spherodised graphite in Micrograph 394, parts of the secondary and tertiary precipitates in the Nickel superalloy.
S6.3 Classifier ablation
Figure S12 shows the result of changing the classifier used for weakly supervised segmentation for different features. When using linear models (Ridge regression, logistic regression) the classical features struggle to differentiate the edges of the primary precipitates from the tertiary precipitates. This is because the classical features are linear, and cannot model non-linear relationships - Weka uses a (non-linear) random forest classifier to overcome this23.
The upsampled DINOv2-S-14 (stride 4, shift 1, 2, flips) features are not sufficiently high resolution to distinguish the precipitates regardless of classifier, though increasing classifier complexity (i.e, using an MLP) does lead to an improved segmentation of the primary vs secondary precipitates.
Adding the classical features in the ‘Hybrid’ scheme allows a good segmentation of each phase, which is not affected by classifier choice, even with linear models. Interestingly, the ‘Hybrid’ scheme seems to perform better (i.e, capture more tertiary precipitates) when using logistic regression or an MLP - we attribute this to smoother decision boundaries. The linear classifiers used default parameters, with 3000 iterations where applicable. The Random Forest was ‘Weka-style’, with 200 trees, 2 features per node and unlimited depth. The MLP had three hidden layers of size 100, ReLU activations and Adam optimizer.
Figure S12: Segmentations of a Nickel superalloy 62 for a fixed set of sparse labels as a function of features and classifier. We see the choice of classifier has little impact on the performance of the hybrid (classical + upsampled DINOv2-S-14) features, but does affect the classical features.
S7 Transformations
Based on experiments and shown in Figure S13, for a ViT model with stride $S$, we found the best range of distances for the shift transforms to be $[1,S/2]$. When combined with the flip transforms this gives $4\text{ (flips)}\times 8\text{ (directions)}\times S/2=32S\text{ transformations}$. Usually $S=4$ so the number of transforms was 64. This could be reduced by using a 4-neighbourhood rather than an 8-neighbourhood, at the cost of slightly reduced performance. Although 4 flip transforms were used (no flip, flip vertical, no flip, flip horizontal), three would also work (no flip, flip vertical, flip horizontal).
@@
Table S1: mIoU across the cell TEM dataset for ViT features (stride 4, no CRF) without and with shift transforms applied - the performance gain is modest for DINOv2 and more marked for DINO.
@@
Table S2: IoU on DUTs foreground object segmentation as a function of transforms applied
Method VOC07 VOC12
DINO 14, 17 0.458 0.462
Ours (DINOv2-S-14, NU) 0.518 0.553
Ours (DINOv2-S-14, NU, multi) 0.614 0.617
Ours (DINOv2-S-14) 0.554 0.572
Ours (DINOv2-S-14, multi) 0.718 0.725
Table S3: CorLoc on VOC07 & VOC12 for our object localization framework. ‘NU’ refers to ‘No Upsampling’: no changes to model stride or pixel shift transforms and nearest-neighbour upsampling of the features. We see that adding our upsampling increases performance.
@@
S7.1 Impact of transformations on WSS predictions
Figure S14: Impact of changing the model stride and adding shift transforms on the resulting weakly-supervised segmentation. Note classical features have been added for all examples, i.e, this is the hybrid scheme. At high strides (stride = model patch size), we get blocky predictions, which is improved as the stride reduces and transformations (shift + flips) are added.
S7.2 Impact of transformations on time/memory cost
Each transformation (a single shift or flip) corresponds to one additional forward pass of the original ViT model (in most cases DINOv2). This means the peak time and memory cost scales linearly with the number of transforms. If they all fit in memory, this can be done in a single forward pass, so the time cost is constant. We call this approach ‘batched’ and the opposite - computing the features of each transformation sequentially - ‘sequential’.
In order to compose two sets of transformations (like flip and pixel-shifts when averaging out positional bias), $T_{1}=\{flip_{LR},flip_{UD}\}$ and $T_{2}=\{shift_{\uparrow,1},shift_{\rightarrow,1},$...\} the total number of transforms is the cartesian product, i.e $N_{transforms}=T_{1}\times T_{2}=|T_{1}||T_{2}|$. Note that for each example there is also a forward pass of the untransformed image. We present the peak time and memory cost with different transformations in Table S4 for DINOv2-S-14 and DINO-S-8 stride 4, noting that it aligns well the findings in Figure 8.
Transformations $N_{t}$ Memory (MB) Time (s)
Batched (DINOv2-S-14)
No transforms 0 281 0.04
Von Neumann shift (1) 4 506 0.12
Moore shift (1) 8 551 0.22
Moore Shift (1, 2) 16 1002 0.42
Moore Shift (1, 2) + flips 64 4008 1.57
Sequential (DINOv2-S-14)
No transforms 0 281 0.04
Von Neumann shift (1) 4 464 0.14
Moore shift (1) 8 466 0.26
Moore Shift (1, 2) 16 472 0.49
Moore Shift (1, 2) + flips 32 508 1.98
Batched (DINO-S-8)
No transforms 0 282 0.03
Von Neumann shift (1) 4 507 0.13
Moore shift (1) 8 553 0.22
Moore Shift (1, 2) 16 1025 0.40
Moore Shift (1, 2) + flips 64 4100 1.55
Table S4: Time and memory cost scaling as a function of added transformations for different DINO models with a stride of 4.